OpenStudy (meilendurcer):

Generate the first 5 terms of this sequence: f(1) = 1 and f(2) = 2, f(n) = f(n - 1) + f(n - 2), for n > 2.

OpenStudy (anonymous):

1,2,3,5,8,13,21.... Is it enough??

OpenStudy (meilendurcer):

yes thanks! can you help with more?

OpenStudy (anonymous):

k...sure..

OpenStudy (meilendurcer):

For the function f(x) = x + 4, what is the ordered pair for the point on the graph when x = 3p?

OpenStudy (anonymous):

f(x)= x + 4 i.e y= = x + 4 when x=3p, y=3p+4 Therefore the ordered pair is (3p,3p+4) Is it ok??

OpenStudy (meilendurcer):

yeah thanks i got it. i'm suppose to fill in the correct number from the graph but i'm not good at reading graphs. f(0) = _______

OpenStudy (anonymous):

I will help you to find it for funtion f(x)=x+4 Now to find f(a) you just have to replace all x in the equation by a.. so now try finding f(0)....

OpenStudy (meilendurcer):

sorry my computer suddenly died

OpenStudy (anonymous):

Sorry even I am going offline...If needed I will be back to help you...Bye..

OpenStudy (meilendurcer):

thanks for the help bye :)

OpenStudy (anonymous):

Did you get the answer??